OpenStudy (anonymous):

y=(2^(x) +1)/ (2^(x) -1)... solve for x. is x=log_2 ((1+y)/(2-y))?

OpenStudy (anonymous):

your equation can become: y(2^x -1) = 2^x + 1 2^x(y-1) = y + 1 2^x = (y+1)/(y-1) ln(2^x) = ln[(y+1)/(y-1)] x = Log_2 [(y+1)/(y-1)]

OpenStudy (anonymous):

oh! my bad about the 2-y! it's supposed to be y-1! thank you again!:D

OpenStudy (anonymous):

yw;)